Meghan has 100 dollars more than Andrew has. After Meghan spends 20 dollars on groceries, she will have five times as much money as Andrew. How much money does Andrew have?
Answer by mananth(16946)

let Andrew have $ x
Megan has x+100
..
x+100-20= 5x
x+80=5x
5x-x =80
4x=80
/4
x= 20
Andrew has $20
..
m.ananth@hotmail.ca